OpenStudy (saifoo.khan):

write each phrase as an algebraic expression 200 less then half of a number

OpenStudy (anonymous):

y = (x/2) - 500. understand it though :)

OpenStudy (saifoo.khan):

ty.

OpenStudy (anonymous):

y = 1/2n - 200

OpenStudy (anonymous):

1/2x-200=y

OpenStudy (anonymous):

you guys should use brackets. they both can look wrong :P

OpenStudy (saifoo.khan):

Everyone's copying Euler's answer. =(

OpenStudy (anonymous):

200-1/2n

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

order in which u put ur #s matters

OpenStudy (anonymous):

or is it 200>1/2n?

OpenStudy (anonymous):

< sorry

OpenStudy (saifoo.khan):

no worries.! ^_^

OpenStudy (anonymous):

yup :)